The payroll of Swifty Company for September 2016 is as follows. Total payroll was $446,000, of which $106,000 is exempt from Social Security tax because it represented amounts paid in excess of $118,500 to certain employees. The amount paid to employees in excess of $7,000 was $400,000. Income taxes in the amount of $78,700 were withheld, as was $8,900 in union dues. The state unemployment tax is 3.5%, but Swifty Company is allowed a credit of 2.3% by the state for its unemployment experience. Also, assume that the current FICA tax is 7.65% on an employee’s wages to $118,500 and 1.45% in excess of $118,500. No employee for Swifty makes more than $125,000. The federal unemployment tax rate is 0.8% after state credit. Prepare the necessary journal entries if the wages and salaries paid and the employer payroll taxes are recorded separately. (Round answers to 0 decimal places, e.g. 5,275. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts .Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1
a
Debit Credit
Salaries and Wages Expense 446000
      Withholding Income Taxes Payable 78700
       FICA Taxes Payable 27547 =(446000-106000)*7.65%+(106000*1.45%)
      Union Dues Payable 8900
      Cash 330853
b
Payroll Tax Expense 28467
       FICA Taxes Payable 27547
       FUTA Taxes Payable 368 =(446000-400000)*0.8%
       SUTA Taxes Payable 552 =(446000-400000)*(3.5%-2.3%)